OpenStudy (anonymous):

Della went to the store to buy 3 pens and 6 notebooks. She spent $20.25. Felipe went to the same store and bought 4 pens and 5 notebooks. He spent $18.75. How much did 1 pen cost?

OpenStudy (anonymous):

@texaschic101 @Gabylovesyou

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (pinklion23):

@Igreen @AnswerMyQuestions @sleepyjess @inowalst

OpenStudy (michele_laino):

If I call with x the cost of 1 pen, and with y the cost of 1 notebook, then the problem is modeled by the subsequent system: \[\large \left\{ \begin{gathered} 3x + 6y = 20.25 \hfill \\ 4x + 5y = 18.75 \hfill \\ \end{gathered} \right.\] please solve it for x

OpenStudy (anonymous):

thank you @Michele_Laino

OpenStudy (anonymous):

@Michele_Laino i still dont get it lol

OpenStudy (michele_laino):

hint: you have to solve that system, using, for example the substitution method, namely, you can find y from the first equation, and then substitute that expression for y, into the second equation. So you can solve that new second equation for x. For example, we have: \[y = \frac{{20.25 - 3x}}{6}\] then I substitute into the second equation: \[4x + 5 \cdot \frac{{20.25 - 3x}}{6} = 18.75\] the please, solve that last equation, for x, and you will get the right answer.

OpenStudy (michele_laino):

@kylefoster58